# Nuclear state and level densities of actinides with the shell-model Monte Carlo

### v1

@@ Appendix A End Matter
Results for other actinides
Figure 4: Same as in Fig. 1, but for other actinides.
Here, we present results for nine other actinides for which Oslo method experiments have not yet been performed. The NSDs, NLDs, and spin distributions at the respective neutron separation energies have been calculated for these nuclei using the same methods as those presented in the main text.
Figures 4 and 5 show the NSDs and NLDs, respectively, for these nuclei. The NSDs of these nuclei show similarly good agreement with the BBF densities as the actinides in the main text. The NLDs and spin-cutoff parameters at the neutron separation energies have been used to compute values of $D_{0}$, shown in Table 2.
Figure 5: Same as in Fig. 3, but for other actinides.
@@
  Exp.     SMMC   Exp.     SMMC
234U 6.845 5/2 47.0 $\pm$ 8.9 49.4 $\pm$ 5.6 8.64 $\pm$ 0.13 0.52 $\pm$ 0.02 0.54 $\pm$ 0.11
235U 5.298 0 11.8 $\pm$ 2.5 6.74 $\pm$ 0.74 8.11 $\pm$ 0.15 $11.2\pm 0.8$ 19.7 $\pm$ 2.3
236U 6.546 7/2 42.1 $\pm$ 7.9 38.3 $\pm$ 3.2 8.56 $\pm$ 0.12 0.45 $\pm$ 0.03 0.53 $\pm$ 0.07
241Pu 5.242 0 $10.3\pm 2.1$ 4.48 $\pm$ 0.57 8.12 $\pm$ 0.18 $13.0\pm 0.5$ 19.7 $\pm$ 4.0
242Pu 6.310 5/2 26.5 $\pm$ 5.4 25.9 $\pm$ 5.3 8.61 $\pm$ 0.17 0.93 $\pm$ 0.08 1.02 $\pm$ 0.23
246Cm 6.458 7/2 24.8 $\pm$ 4.6 26.7 $\pm$ 4.3 8.60 $\pm$ 0.18 0.79 $\pm$ 0.05 0.78 $\pm$ 0.18
247Cm 5.156 0 $4.76\pm 1.23$ 3.29 $\pm$ 0.42 8.02 $\pm$ 0.20 30.0 $\pm$ 5.0 39.4 $\pm$ 5.4
248Cm 6.213 9/2 14.0 $\pm$ 4.2 19.8 $\pm$ 2.4 8.47 $\pm$ 0.18 1.2 $\pm$ 0.3 0.87 $\pm$ 0.11
250Cf 6.624 9/2 23.0 $\pm$ 4.9 41.1 $\pm$ 4.9 8.66 $\pm$ 0.16 0.76 $\pm$ 0.10 0.43 $\pm$ 0.06
Table 2: Same as in Table 1, but for other actinides

### v2

@@ Appendix A End matter
Results for other actinides.— In the following we present results for nine other actinides for which Oslo method experiments have not been performed but for which neutron resonance data and low-lying level counting data exist. The NSDs, NLDs, and spin distributions at the respective neutron separation energies have been calculated for these nuclei using the same methods as those presented in the main text.
Figures 4 and 5 show the NSDs and NLDs, respectively, for these nuclei. The NSDs of these nuclei show similarly good agreement with the BBF densities as the actinides in the main text. Table 2 provides the corresponding experimental and calculated values of the NLDs, spin-cutoff parameters and $D_{0}$ at the neutron separation energies.
Figure 4: As in Fig. 1, but for other actinides.
Figure 5: As in Fig. 3, but for other actinides.
@@
  Exp.     SMMC rigid body    SMMC   Exp.     SMMC
234U 6.845 5/2 46.8 $\pm$ 8.9 49.4 $\pm$ 5.6 rigid 8.24 rigid 8.64 $\pm$ 0.13 0.52 $\pm$ 0.02 0.54 $\pm$ 0.11
235U 5.298 0 11.7 $\pm$ 2.5 6.74 $\pm$ 0.74 8.07 8.11 $\pm$ 0.15 $11.2\pm 0.8$ 19.7 $\pm$ 2.3
236U 6.546 7/2 41.9 $\pm$ 7.8 38.3 $\pm$ 3.2 8.15 8.56 $\pm$ 0.12 0.45 $\pm$ 0.03 0.53 $\pm$ 0.07
241Pu 5.242 0 $10.2\pm 2.1$ 4.48 $\pm$ 0.57 8.12 8.12 $\pm$ 0.18 $13.0\pm 0.5$ 19.7 $\pm$ 4.0
242Pu 6.310 5/2 26.3 $\pm$ 5.4 25.9 $\pm$ 5.3 8.27 8.61 $\pm$ 0.17 0.93 $\pm$ 0.08 1.02 $\pm$ 0.23
246Cm 6.458 7/2 24.6 $\pm$ 4.6 26.7 $\pm$ 4.3 8.30 8.60 $\pm$ 0.18 0.79 $\pm$ 0.05 0.78 $\pm$ 0.18
247Cm 5.156 0 $4.71\pm 1.22$ 3.29 $\pm$ 0.42 8.37 8.02 $\pm$ 0.20 30.0 $\pm$ 5.0 39.4 $\pm$ 5.4
248Cm 6.213 9/2 13.9 $\pm$ 4.1 19.8 $\pm$ 2.4 8.32 8.47 $\pm$ 0.18 1.2 $\pm$ 0.3 0.87 $\pm$ 0.11
250Cf 6.624 9/2 22.7 $\pm$ 4.8 41.1 $\pm$ 4.9 8.51 8.66 $\pm$ 0.16 0.76 $\pm$ 0.10 0.43 $\pm$ 0.06
Table 2: As in Table 1, but for other actinides